# Double Higgs Production via Vector Boson Fusion in SMEFT

### v1

@@ Abstract
While gluon fusion dominates Higgs pair production at the LHC, vector boson fusion (VBF) offers a unique window into Beyond the Standard Model (BSM) physics through its distinctive kinematic features and direct sensitivity to Higgs-vector boson interactions. We perform a comprehensive analysis of double Higgs production via VBF in the Standard Model Effective Field Theory (SMEFT), systematically investigating how dimension-6 and dimension-8 bosonic operators – particularly those involving field derivatives – can enhance the production rate. We identify the most relevant Wilson coefficients (WCs) affecting the trilinear Higgs coupling ($hhh$) and Higgs-vector boson interactions ($hVV$, $hhVV$). Using constraints from global fits and interpolating fit results for unconstrained WCs with Naive Dimensional Analysis, we assess their effects on the $VV\to hh$ ($V=W,Z$) scattering amplitudes and cross-sections. Our analysis includes a study of EFT convergence and validity in models with scalar extensions of the SM. Numerical simulations for the planned High-Luminosity LHC experiment (HL-LHC) in general reveal only a modest and challenging to detect enhancement of the VBF di-Higgs production rate over the SM prediction. However, we show that in optimistic scenarios, such a process could be observed at the HL-LHC. In certain cases, when the enhancement is dominated by the dimension-6 or dimension-8 operators containing field derivatives (and thus leading to stronger energy-dependent effects), this channel becomes competitive with di-Higgs production via gluon fusion. This work highlights the role of VBF di-Higgs production as a complementary channel for probing anomalous Higgs couplings and their impact on BSM physics.
@@ 1 Introduction
The two main production channels for Higgs boson pair production in the SM are gluon fusion and vector boson fusion processes [8, 9, 10, 11]. While gluon fusion $(gg\to hh)$ [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] dominates in the SM, vector boson fusion $(qq\to qqVV\to qqhh)$ [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] constitutes a complementary channel with distinct experimental signatures, including presence of forward jets and low hadronic activity in the central region. Moreover, although VBF di-Higgs production is suppressed by an order of magnitude relative to ggF channel, it is highly sensitive to anomalous Higgs couplings - namely, the trilinear coupling $hhh$ and Higgs-gauge boson couplings $hVV$ and $hhVV$ (where “$V$” stands for either a $W$-boson or a $Z$-boson). Such couplings commonly arise in various BSM theories. For reference, the leading-order tree-level diagrams contributing to di-Higgs VBF production at LHC are shown in Figure 1, with anomalous couplings marked by red blobs.
@@
The Standard Model Effective Field Theory (SMEFT) provides a model-independent framework to parameterise BSM effects through higher-dimensional operators, suppressed by a cutoff scale $\Lambda$. These operators modify Higgs and gauge boson interactions, leading to deviations in cross-sections and kinematic distributions. In the context of double Higgs VBF production, dimension-6 and dimension-8 operators can, in principle, significantly alter the scattering amplitudes, offering a window into new physics entering the trilinear Higgs coupling that may be inaccessible in single-Higgs processes.
Over the years, numerous studies have investigated double Higgs production in the context of various BSM scenarios [59, 60, 61, 62, 63, 64, 65, 66]. A number of recent analyses have employed EFT approach, including both SMEFT and HEFT [67], to study single, double, and multi Higgs production. Some focused on the ggF channel [68, 69, 70, 71, 72, 73, 74, 75, 76], while others examined modifications to the trilinear Higgs coupling [77, 78, 79, 80, 81, 82, 83], or concentrated on the VBF channel [84, 85, 86, 87, 88, 89, 90, 91].
@@
The paper is organised as follows. Section 2 identifies the dimension-6 and dimension-8 SMEFT operators relevant for this analysis, and outlines our assumptions leading to bounds on maximal values of their Wilson coefficients used further in numerical analysis. In Section 3 we discuss the EFT convergence and accuracy, illustrating it with examples of the three BSM models involving heavy scalar singlets and triplets, by comparing the amplitudes calculated in renormalizable UV-theories with the corresponding ones obtained in the decoupling limit. Analytical formulae for $VV\to hh$ helicity amplitudes and cross-sections are presented in Section 4 and Appendices A and B. Finally, Section 5 contains the results of our numerical simulations, directly addressing the question of maximal potential enhancement of VBF di-Higgs production in SMEFT for LHC and HL-LHC experiments. Section 6 concludes the paper.
@@ 2.1 Relevant operators
To affect the process of double Higgs production via VBF, the SMEFT operators can be inserted into three types of vertices: $hhh$, $hVV$, and $hhVV$ (red dots on diagrams in Figure 1, $h$ indicates the Higgs and $V$ indicates the gauge boson). The relevant classes of the SMEFT operators that can impact this process are all summarised in Table 1.
@@
$hhh$ $\varphi^{6}$, $\varphi^{4}D^{2}$ $\varphi^{8}$, $\varphi^{6}D^{2}$
$hVV$ $\varphi^{4}D^{2}$, $X^{2}\varphi^{2}$ $\varphi^{6}D^{2}$, $X^{2}\varphi^{4}$, $X\varphi^{4}D^{2}$
$hhVV$ $\varphi^{4}D^{2}$, $X^{2}\varphi^{2}$ $\varphi^{4}D^{4}$, $\varphi^{6}D^{2}$, $X^{2}\varphi^{4}$, $X^{2}\varphi^{2}D^{2}$
@@
Some important comments are in place here. As we are interested in the most general features of the VBF part of the process, we neglect the impact of the SMEFT operators on the external quark lines and on further decays of the Higgs boson pair into the final states. At the dimension-6 level, the primary potential contributions may originate from the dipole operators $Q_{uW}^{11},Q_{dW}^{11}$ that are chirality-flipping, as well as the chiral operators $Q_{\varphi ud}^{11}$ and $[Q_{\varphi q}^{(3)}]^{11}$. Especially the latter could amplify the longitudinal $W$-boson scattering, and as it was shown in ref. [92] for $WW\to WW$ scattering, these operators result in distinct kinematic distributions, possibly leading to the same effects in $WW\to hh$. However, they are strongly constrained by the $Z$-pole precision observables, e.g. for $[Q_{\varphi q}^{(3)}]^{11}$, a recent electroweak analysis [93] shows a tight bound on its corresponding WC, of less than $0.02\,\frac{1}{\mathrm{TeV}^{2}}$. Despite that, the concern about such fermionic operators affecting studied process remains, especially at the dimension-8 level. Effects of a subset of such operators have been recently studied in Ref. [76] in the case of single Higgs production. Additionally, we neglect all CP-violating operators. We also do not consider the effects of the Renormalization Group Equation (RGE) running of the WCs, although we conducted a brief examination of their magnitude, finding them negligible.
@@ 2.2 Bounds on Wilson coefficients
[…]
However, within linear field redefinitions in SMEFT, the trilinear Higgs coupling $hhh$ is a momentum-dependent quantity through the WCs $C_{\varphi D}$ and $C_{\varphi\Box}$ and related WCs associated with dimension-8 operators. Under the assumption that all other WCs are negligibly small compared to $C_{\varphi}$ and $C_{\varphi^{8}}$, this limit can be translated into a bound on the following combination in the $\{G_{F},m_{W},m_{Z}\}$-input parameter scheme [99, 75],
@@
$-\frac{13.8}{\text{TeV}^{2}}\lesssim C_{\varphi}\lesssim\frac{4.7}{\text{TeV}^{2}}\;.$ (2.3)
While our study adopts conservative bounds on $C_{\varphi}$, we note that recent trilinear coupling analysis [100] suggests that somewhat stronger constraints may apply.
@@ 3.1 Real scalar model
The matching onto the SMEFT of a heavy scalar singlet model has been worked out in several studies, with dimension-6 operators at tree-level [105, 106] and one-loop [107, 108] as well as dimension-8 operators [109]. By incorporating all renormalizable interactions between the heavy singlet $S$ and the SM fields, we have
@@
where $\mathcal{L}_{SM}$ is the SM-Lagrangian in the notation of Refs. [110, 104] and $\varphi$ is the SM Higgs doublet. Besides the heavy singlet mass $M$, all other parameters are dimensionless. By promoting the parameters of the theory into auxiliary fields, a background $Z_{2}$-symmetry under which $\varphi\to\varphi$ and $S\to-S$, assigns charges $(-1)$ to $\alpha$ and $\beta$ and $(+1)$ to all other couplings. The Wilson coefficients in front of the $Z_{2}$-invariant EFT expansions in terms of $v^{2}/M^{2}$ (or $s/M^{2}$ when calculating momentum-dependent terms in scattering amplitudes) should also have an $Z_{2}$-invariant polynomial form,
@@
where $\lambda$ is the quartic Higgs coupling in $\mathcal{L}_{SM}$: $-\frac{\lambda}{2}(\varphi^{\dagger}\varphi)^{2}$.
@@
so that $\alpha\lesssim 4\pi$ and maximal value of $C_{\varphi\Box}$ can saturate the NDA limit. Therefore, the SM extension with a real scalar provides a simple, explicit example of a model that can really produce large values for WCs in the low-energy limit555 However, one should note that some WCs may be constrained by additional conditions, such as the properties of the Higgs potential. Higher order operators contribute to it, ultimately affecting its stability and positions of minima. We do not study these issues in detail here, as our goal is simply to provide an example of consistent UV models which may produce large WCs..
@@
To address the question of EFT convergence, let us note that comparison of Eq. (3.3) with Eqs. (3.4) and (3.5) shows that WCs of all higher order operators contributing to the dominant terms in the amplitude are proportional to $\alpha^{2}$ and not to the higher powers of $\alpha$. Consequently, they do not grow uncontrollably with increasing mass dimension and do not endanger the EFT series convergence. This remark is also in agreement with NDA as on dimensional ground all such couplings should be proportional to $g_{*}^{2}$. From Eq. (3.3) it is clear that EFT series always converges if we are below the heavy mass threshold, $s<M^{2}$, independently of WC values. Other WCs which may contribute to sub-dominant terms in the amplitude are proportional to higher powers of UV parameters and thus can be larger, but they are always multiplied by powers of $(v/M)$ and therefore should not affect EFT convergence as long as condition (2.7) is fulfilled.
@@ 3.2 Triplet scalar models
As can be seen from Eq. (3.3), the dominant contribution to $W_{L}W_{L}\to hh$ and $Z_{L}Z_{L}\to hh$ in real scalar model has always opposite sign to SM term $-\lambda_{eff}$ ($\lambda_{eff}$ should be positive for Higgs potential to be stable and to have proper minima). Thus, in such a model both processes are suppressed compared to the SM, unless the $\alpha$-parameter is huge and BSM terms dominate the SM amplitude. This may not be the case for scalar triplet extensions.
@@
[…]
@@
In summary, we demonstrated with explicit examples that it is possible to find renormalizable UV-models which produce at low energies large WCs in SMEFT, that may increase or decrease the di-Higgs production rates through VBF and are consistent with EFT expansion convergence.
@@ 3.3 EFT accuracy for large $\sqrt{s}$
Figure 3: The plots compare the relative accuracy of the $W_{L}W_{L}\to hh$ high energy scattering cross-section as predicted by the SMEFT and underlying UV models. The dashed lines represent $\Delta_{EFT}$ from Eq. (3.27) as a function of $\sqrt{s}/M$. The orange curve shows the EFT result without dimension-8 or higher terms, while the blue and red curves include dimension-8 terms in the amplitude, either truncated or fully kept in its square, respectively. The dotted lines mark the ±$\pm 10\%$ range of relative accuracy. The solid black line shows the deviation of the cross-section ratios from unity in the scalar models relative to the SM. The upper panels present the findings for a real scalar model, and the lower panels show the findings for a model extended with both neutral and charged triplets.
@@
Following the above discussion, in inequality of Eq. (3.28) one should rather set $x\approx 0.7$. However, we use the more relaxed value $x=1$ because as we shall see in Section 5, even such condition in most cases leaves little room for process enhancement, meaning that the more stringent bound would lead to even less optimistic predictions.
It’s also important to note that using only linear terms in dimension-6 Wilson coefficients yields considerably less accurate results. Furthermore, Figure 3 demonstrates that when the discrepancy between EFT and UV predictions becomes significant for stronger couplings, the EFT better reproduces the full result if the amplitude is truncated at dimension-8 terms, even when formally higher-order terms (up to dimension-12) arising from squaring the amplitude are retained in the cross-section. This conclusion aligns with the analysis in Ref. [116], which explicitly compares UV model results with their corresponding EFT limits and finds that truncating EFT amplitudes alone leads to better agreement with the full theory. We adopt this same approach in our paper and give a few more details in the following subsection.

### v2

@@ Abstract
While gluon fusion dominates Higgs pair production at the LHC, vector boson fusion (VBF) offers a unique window into Beyond the Standard Model (BSM) physics through its distinctive kinematic features and direct sensitivity to Higgs-vector boson interactions. We perform a comprehensive analysis of double Higgs production via VBF in the Standard Model Effective Field Theory (SMEFT), systematically investigating how dimension-6 and dimension-8 bosonic operators – particularly those involving field derivatives – can enhance the production rate. We identify the most relevant Wilson coefficients (WCs) affecting the trilinear Higgs coupling (hhhhhh) and Higgs-vector boson interactions ($hVV$, hhVVhhVV). Using constraints from global fits and interpolating fit results for unconstrained WCs with Naive Dimensional Analysis, we assess their effects on the $VV\to hh$ ($V=W,Z$) scattering amplitudes and cross-sections. Our analysis includes a study of EFT convergence and validity in models with scalar extensions of the SM. Numerical simulations for the planned High-Luminosity LHC experiment (HL-LHC) in general reveal only a modest and challenging to detect enhancement of the VBF di-Higgs production rate over the SM prediction. However, we show that in optimistic scenarios, such a process could be observed at the HL-LHC. In certain cases, when the enhancement is dominated by the dimension-6 or dimension-8 operators containing field derivatives (and thus leading to stronger energy-dependent effects), this channel becomes competitive with di-Higgs production via gluon fusion. This work highlights the role of VBF di-Higgs production as a complementary channel for probing anomalous Higgs couplings and their impact on BSM physics.
@@ 1 Introduction
The two main production channels for Higgs boson pair production in the SM are gluon fusion and vector boson fusion processes [8, 9, 10, 11]. While gluon fusion $(gg\to hh)$ [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] dominates in the SM, vector boson fusion $(qq\to qqVV\to qqhh)$ [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] constitutes a complementary channel with distinct experimental signatures, including presence of forward jets and low hadronic activity in the central region. Moreover, although VBF di-Higgs production is suppressed by an order of magnitude relative to ggF channel, it is highly sensitive to anomalous Higgs couplings - namely, the trilinear coupling hhhhhh and Higgs-gauge boson couplings $hVV$ and $hhVV$ (where “VV” stands for either a $W$-boson or a $Z$-boson). Such couplings commonly arise in various BSM theories. For reference, the leading-order tree-level diagrams contributing to di-Higgs VBF production at LHC are shown in Figure 1, with anomalous couplings marked by red blobs.
@@
The Standard Model Effective Field Theory (SMEFT) [59, 60, 61, 62, 63] provides a model-independent framework to parameterise BSM effects through higher-dimensional operators, suppressed by a cutoff scale $\Lambda$. These operators modify Higgs and gauge boson interactions, leading to deviations in cross-sections and kinematic distributions. In the context of double Higgs VBF production, dimension-6 and dimension-8 operators can, in principle, significantly alter the scattering amplitudes, offering a window into new physics entering the trilinear Higgs coupling that may be inaccessible in single-Higgs processes.
Over the years, numerous studies have investigated double Higgs production in the context of various BSM scenarios [64, 65, 66, 67, 68, 69, 70, 71]. A number of recent analyses have employed EFT approach, including both SMEFT and Higge Effective Field Theory (HEFT) [72, 73, 74, 75, 76], to study single, double, and multi Higgs production. Some focused on the ggF channel [77, 78, 79, 80, 81, 82, 83, 84, 85], while others examined modifications to the trilinear Higgs coupling [86, 87, 88, 89, 90, 91, 92], or concentrated on the VBF channel [93, 94, 95, 96, 97, 98, 99, 100].
@@
The paper is organised as follows. Section 2 identifies the dimension-6 and dimension-8 SMEFT operators relevant for this analysis, and outlines our assumptions leading to bounds on maximal values of their Wilson coefficients used further in numerical analysis. In Section 3 we discuss the EFT convergence and accuracy, illustrating it with examples of the three BSM models involving heavy scalar singlets and triplets, by comparing the amplitudes calculated in renormalisable UV-theories with the corresponding ones obtained in the decoupling limit. Analytical formulae for $VV\to hh$ helicity amplitudes and cross-sections are presented in Section 4 and Appendices A and B. Finally, Section 5 contains the results of our numerical simulations, directly addressing the question of maximal potential enhancement of VBF di-Higgs production in SMEFT for LHC and HL-LHC experiments. Section 6 concludes the paper.
@@ 2.1 Relevant operators
To affect the process of double Higgs production via VBF, the SMEFT operators can be inserted into three types of vertices: hhhhhh, hVVhVV, and $hhVV$ (red dots on diagrams in Figure 1, $h$ indicates the Higgs and $V$ indicates the gauge boson). The relevant classes of the SMEFT operators that can impact this process are all summarised in Table 1.
@@
hhhhhh $\varphi^{6}$, $\varphi^{4}D^{2}$ $\varphi^{8}$, $\varphi^{6}D^{2}$
hVVhVV $\varphi^{4}D^{2}$, $X^{2}\varphi^{2}$ $\varphi^{6}D^{2}$, $X^{2}\varphi^{4}$, $X\varphi^{4}D^{2}$
hhVVhhVV $\varphi^{4}D^{2}$, $X^{2}\varphi^{2}$ $\varphi^{4}D^{4}$, $\varphi^{6}D^{2}$, $X^{2}\varphi^{4}$, $X^{2}\varphi^{2}D^{2}$
@@
Some important comments are in place here. As we are interested in the most general features of the VBF part of the process, we neglect the impact of the SMEFT operators on the external quark lines and on further decays of the Higgs boson pair into the final states. At the dimension-6 level, the primary potential contributions may originate from the dipole operators $Q_{uW}^{11},Q_{dW}^{11}$ that are chirality-flipping, as well as the chiral operators $Q_{\varphi ud}^{11}$ and $Q_{\varphi q}^{(3)}]^{11}$. Especially the latter could amplify the longitudinal $W$-boson scattering, and as it was shown in ref. [101] for $WW\to WW$ scattering, these operators result in distinct kinematic distributions, possibly leading to the same effects in $WW\to hh$. However, they are strongly constrained by the $Z$-pole precision observables, e.g. for $Q_{\varphi q}^{(3)}]^{11}$, a recent electroweak analysis [102] shows a tight bound on its corresponding WC, of less than $0.02\,\frac{1}{\mathrm{TeV}^{2}}$. Despite that, the concern about such fermionic operators affecting studied process remains, especially at the dimension-8 level. Effects of a subset of such operators have been recently studied in Ref. [85] in the case of single Higgs production. Additionally, we neglect all CP-violating operators. We also do not consider the effects of the Renormalisation Group Equation (RGE) running of the WCs. Although it has been shown in the literature (see, e.g., [80, 103, 104, 105, 106]) that RGE effects are often non-negligible, we have verified that, in our case, they can change the values of the dimension-6 WCs by about $\mathcal{O}(20\%)$. This may slightly modify our results, but with no significant impact on the final conclusions.
@@ 2.2 Bounds on Wilson coefficients
The maximum allowed values for $C_{\varphi}$ (not shown in Table 3) can be obtained from the current allowed 95% confidence level intervals for the trilinear Higgs coupling modifier $\kappa_{\lambda_{3}}=\lambda_{3}/\lambda_{3}^{SM}$. Bounds on $\kappa_{\lambda_{3}}$ derived from combined single- and double Higgs production searches read (with other couplings set to their SM predictions) [^4] There is also new CMS measurement constraining $\kappa_{\lambda_{3}}$ to similar range, $\kappa_{\lambda_{3}}\in[-1.39,\,7.02]$, based on di-Higgs production only [7].:
@@
However, within linear field redefinitions in SMEFT, the trilinear Higgs coupling hhhhhh is a momentum-dependent quantity through the WCs $C_{\varphi D}$ and $C_{\varphi\Box}$ and related WCs associated with dimension-8 operators. Under the assumption that all other WCs are negligibly small compared to $C_{\varphi}$ and $C_{\varphi^{8}}$, this limit can be translated into a bound on the following combination in the $\{G_{F},m_{W},m_{Z}\}$-input parameter scheme [112, 84],
@@
$\frac{13.8}{\text{TeV}^{2}}\lesssim C_{\varphi}\lesssim\frac{4.7}{\text{TeV}^{2}}\;.$ (2.3)
While our study adopts conservative bounds on $C_{\varphi}$, we note that recent trilinear coupling analysis [113, 114] suggests that somewhat stronger constraints may apply.
@@ 3.1 Real scalar model
The matching onto the SMEFT of a heavy scalar singlet model has been worked out in several studies, with dimension-6 operators at tree-level [119, 120] and one-loop [121, 122] as well as dimension-8 operators [123]. By incorporating all renormalisable interactions between the heavy singlet $S$ and the SM fields, we have
@@
where $\mathcal{L}_{SM}$ is the SM-Lagrangian in the notation of Refs. [63, 118] and $\varphi$ is the SM Higgs doublet. Besides the heavy singlet mass $M$, all other parameters are dimensionless. By promoting the parameters of the theory into auxiliary fields, a background $Z_{2}$-symmetry under which $\varphi\to\varphi$ and $S\to-S$, assigns charges (1)(-1) to $\alpha$ and $\beta$ and $(+1)$ to all other couplings. The Wilson coefficients in front of the $Z_{2}$-invariant EFT expansions in terms of $v^{2}/M^{2}$ (or $s/M^{2}$ when calculating momentum-dependent terms in scattering amplitudes) should also have an $Z_{2}$-invariant polynomial form,
@@
where $\lambda$ is the quartic Higgs coupling in $\mathcal{L}_{SM}$: $\frac{\lambda}{2}(\varphi^{\dagger}\varphi)^{2}$.
@@
so that $\alpha\lesssim 4\pi$ and maximal value of $C_{\varphi\Box}$ can saturate the NDA limit. Therefore, the SM extension with a real scalar provides a simple, explicit example of a model that can really produce large values for WCs in the low-energy limit [^5] However, one should note that some WCs may be constrained by additional conditions, such as the properties of the Higgs potential. Higher order operators contribute to it, ultimately affecting its stability and positions of minima. We do not study these issues in detail here, as our goal is simply to provide an example of consistent UV models which may produce large WCs..
@@
To address the question of EFT convergence, let us note that comparison of Eq. (3.3) with Eqs. (3.4) and (3.5) shows that WCs of all higher order operators contributing to the dominant terms in the amplitude are proportional to $\alpha^{2}$ and not to the higher powers of $\alpha$. Consequently, they do not grow uncontrollably with increasing mass dimension and do not endanger the EFT series convergence. This remark is also in agreement with NDA as on dimensional ground all such couplings should be proportional to $g_{*}^{2}$. From Eq. (3.3), and, at tree level in $\hbar$-expansion, it is clear that EFT series always converges if we are below the heavy mass threshold, $s<M^{2}$, independently of WC values. Other WCs which may contribute to sub-dominant terms in the amplitude are proportional to higher powers of UV parameters and thus can be larger, but they are always multiplied by powers of $(v/M)$ and therefore should not affect EFT convergence as long as condition (2.7) is fulfilled.
@@ 3.2 Triplet scalar models
As can be seen from Eq. (3.3), the dominant contribution to $W_{L}W_{L}\to hh$ and $Z_{L}Z_{L}\to hh$ in real scalar model has always opposite sign to SM term $\lambda_{eff}$ ($\lambda_{eff}$ should be positive for Higgs potential to be stable and to have proper minima). Thus, in such a model both processes are suppressed compared to the SM, unless the $\alpha$-parameter is huge and BSM terms dominate the SM amplitude. This may not be the case for scalar triplet extensions.
@@
$\mathcal{M}_{00}^{G^{0}G^{0}\to hh}(s)$ $=-\lambda\ +\ 2\,\alpha_{T_{2}}^{2}\,M^{2}\,\biggl[\frac{1}{s-M^{2}}-\frac{1}{t-M^{2}}-\frac{1}{u-M^{2}}\biggr]$
@@
In summary, we demonstrated with explicit examples that it is possible to find renormalisable UV-models which produce at low energies large WCs in SMEFT, that may increase or decrease the di-Higgs production rates through VBF and are consistent with EFT expansion convergence.
@@ 3.3 EFT accuracy for large $\sqrt{s}$
Figure 3: The plots compare the relative accuracy of the $W_{L}W_{L}\to hh$ high energy scattering cross-section as predicted by the SMEFT and underlying UV models. The dashed lines represent $\Delta_{EFT}$ from Eq. (3.27) as a function of $\sqrt{s}/M$. The orange curve shows the EFT result with dimension-6 terms in the amplitude alone, with higher-order terms truncated in the cross-section. The blue and red curves include dimension-8 terms in the amplitude, either truncated or fully kept in its square, respectively. The dotted lines mark the $\pm 10\%$ range of relative accuracy. The solid black line shows the deviation of the cross-section ratios from unity in the scalar models relative to the SM. The upper panels present the findings for a real scalar model, and the lower panels show the findings for a model extended with both neutral and charged triplets.
@@
Although the value $x\approx 0.7$ is theoretically preferable for inequality (3.28), we adopt the more relaxed value $x=1$. This choice is justified by our numerical analysis in Section 5, which shows that even $x=1$ severely limits potential process enhancements. Consequently, the stricter bound $x\approx 0.7$ would yield even less favorable predictions.
It is also important to note that retaining only linear terms in dimension-6 Wilson coefficients yields considerably less accurate results. Furthermore, Figure 3 demonstrates that when the discrepancy between EFT and UV predictions becomes significant for stronger couplings. In some cases, EFT more accurately reproduces the full result if the amplitude is truncated at dimension-8 terms, even when formally higher-order terms (up to dimension-12) arising from squaring the amplitude are retained in the cross-section. More details are provided in the following subsection.